# RaLiBEV: Radar and LiDAR BEV Fusion Learning for Anchor Box Free Object Detection Systems

### 01 Dec 2023 v4

@@ V Conclusion
In this paper, we introduce RaLiBEV, an innovative method for fusing LiDAR and radar data. According to the experiment results, this method effectively showcases the significant role that radar can play in fusing autonomous driving perception technology, especially in adverse weather. To enhance the performance of the model, especially for high-precision detection, novel label assignment and transformer-based fusion strategies are proposed. Specifically, the GACHIPS strategy is proposed to dynamically adjust the classification and regression positive sample position, to solve the inconsistency problem in the loss procedure. Meanwhile, the DQMITBF is proposed to aggregate more aligned features from different sensor data. Due to the introduction of the dense learnable query in the fusion module, RaLiBEV can fuse the features from radar and LiDAR in a more flexible way. It’s noteworthy that by incorporating the GACHIPS label assignment into a simple fusion architecture, the model can attain performance on par with the more intricate transformer fusion. The experiments with ORR dataset show that RaLiBEV exceeds all existing state-of-the-art approaches by a large margin.

### 06 Feb 2024 v5

@@ V Conclusion
In this paper, we introduce RaLiBEV, an innovative method for fusing LiDAR and radar data. According to the experiment results, this method effectively showcases the significant role that radar can play in fusing autonomous driving perception technology, especially in adverse weather. To enhance the performance of the model, especially for high-precision detection, novel label assignment and transformer-based fusion strategies are proposed. Specifically, the GACHIPS strategy is proposed to dynamically adjust the classification and regression positive sample position, to solve the inconsistency problem in the loss procedure. Meanwhile, the DQMITBF is proposed to aggregate weighted features from different sensor data. Thanks to the interaction of the dense learnable query with radar and LiDAR in the symmetrical manner, RaLiBEV can swtich the focus on different sensor modalities adaptively, according to change of the perception environment and weather conditions. It’s also noteworthy that by incorporating the GACHIPS label assignment into a simple fusion architecture, the model can attain performance on par with the more intricate transformer fusion. The experiments with ORR dataset show that RaLiBEV exceeds all existing state-of-the-art approaches by a large margin.